# Epidemiology-informed Graph Neural Network for Heterogeneity-aware Epidemic Forecasting

### v1

@@ IV-F Ablation Study
  • HeatGNN w/o TG: Remove the TG module and use only the STGL module for direct forecasting, while retaining the EIEL module to guide the output of the framework.
  • HeatGNN w/o TP: Remove the EIEL and TG modules and only keep the STGL module for forecasting.
@@
The ablation study results of RMSE and PCC are shown in Table III. The results show that our proposed HeatGNN performs robustly across most cases, except for horizon=4 in the US-States dataset. The physical loss incorporates domain knowledge of epidemic transmission, particularly benefiting high-variability regions (e.g., Japan-Prefectures and US-Regions) and long-term forecasting. Removing the physics loss results in the model lacking intrinsic epidemic transmission mechanism constraints, which increases the forecasting error and reduces its ability to capture trends. However, its impact is limited in the US-States dataset, which is low volatility and relatively uniform data. The TG module is crucial for modeling cross-regional epidemic transmission. After removing the TG module, the model is significantly weaker at capturing the intrinsic epidemic transmission relationship between locations. While it has a similar role to the physics loss across datasets, its impact on forecasting performance is more pronounced. Additionally, the EIEL module enhances the effectiveness of the TG module, further improving the model’s ability to capture epidemic transmission. Relying solely on the STGL module is insufficient for addressing complex epidemic forecasting tasks. When both the TG and EIEL modules are removed, the model struggles to accurately capture epidemic transmission over extended time horizons, resulting in a significant decline in forecasting performance. These findings suggest that the modules interact synergistically to achieve optimal performance in HeatGNN.
TABLE III: Ablation test results on three datasets.
RMSE($\downarrow$) 2 3 4 5
Japan-Prefectures
HeatGNN w/o PL 1230.134 1360.714 1407.590 1453.051
HeatGNN w/o TG 1228.951 1331.692 1421.312 1438.525
HeatGNN w/o TP 1278.808 1322.376 1441.331 1451.568
HeatGNN 1148.850 1260.150 1295.697 1378.207
US-Regions
HeatGNN w/o PL 631.080 799.733 869.248 942.947
HeatGNN w/o TG 739.933 811.149 854.826 1039.151
HeatGNN w/o TP 726.639 861.203 947.868 977.450
HeatGNN 540.573 693.655 781.513 852.311
US-States
HeatGNN w/o PL 146.414 168.938 172.590 188.129
HeatGNN w/o TG 148.601 176.770 176.644 201.611
HeatGNN w/o TP 152.508 181.273 169.937 206.313
HeatGNN 142.433 164.788 177.624 186.280
PCC($\uparrow$) 2 3 4 5
Japan-Prefectures
HeatGNN w/o PL 0.900 0.889 0.876 0.863
HeatGNN w/o TG 0.891 0.895 0.898 0.871
HeatGNN w/o TP 0.890 0.893 0.893 0.853
HeatGNN 0.917 0.907 0.910 0.884
US-Regions
HeatGNN w/o PL 0.924 0.892 0.875 0.864
HeatGNN w/o TG 0.923 0.891 0.881 0.838
HeatGNN w/o TP 0.927 0.895 0.866 0.854
HeatGNN 0.941 0.908 0.885 0.866
US-States
HeatGNN w/o PL 0.951 0.933 0.929 0.917
HeatGNN w/o TG 0.951 0.927 0.930 0.907
HeatGNN w/o TP 0.951 0.925 0.932 0.902
HeatGNN 0.953 0.939 0.929 0.921
IV-G Hyperparameter Sensitivity Analysis
This section addresses RQ3. We investigate how the forecasting performance varies with some hyperparameters.
IV-G1 Window size
To evaluate whether the model is sensitive to the length of historical data, we evaluate different window sizes ranging from 20 to 60. Figures 3 (a), (b), and (c) show the RMSE and PCC results in different window sizes. For the Japan-Prefectures dataset, forecasting performance improves significantly with larger window sizes, as evidenced by a gradual decrease in RMSE and an increase in PCC. This result suggests that a larger window size provides more comprehensive historical information, enhancing forecasting accuracy. In contrast, model performance remains stable across all window sizes for the US-Regions dataset, which indicates that the model is not sensitive to the window size change in this dataset. For the US-States dataset, the optimal performance is observed at a window size of 60. In summary, while increasing the window size enhances model performance, the improvement becomes limited after a threshold.
IV-G2 Horizon size
To evaluate the sensitivity of the model to the size of the forecasting horizon, we analyze its performance across horizons ranging from 1 to 6. Figures 3 (d), (e), and (f) show the RMSE and PCC results for different horizons. In the Japan-Prefectures dataset, as the forecasting horizon increases, RMSE gradually rises while PCC decreases, indicating a decline in forecasting accuracy with longer horizons. Similar trends are observed in the US-Regions and US-States datasets. Comparing the results across the three datasets, we find that datasets with more significant fluctuations (Japan-Prefectures), shorter forecasting horizons, and more accurate predictions, while datasets with lower variability (US-Regions and US-States) maintain relatively robust forecasting performance even with longer horizons.
IV-G3 Loss weight
To evaluate the sensitivity of the model to the balance weight between physical and data loss, we examine the effect of varying $\lambda$ values from 0 to 1 on forecasting performance. Figures 3 (g), (h), and (i) show the RMSE and PCC results under different $\lambda$. On the Japanese-Prefectures dataset, the PCC of the model remains stable as $\lambda$ increases (i.e., the weight of the physical loss increases), while the RMSE increases slightly at $\lambda$= 0.2. This result indicates that performance is slightly affected by low physical loss weights, and the overall trend remains stable. Comparing the results across the three datasets, we observe that the Japan-Prefectures dataset performs slightly worse under low physical loss weights. In contrast, the US-States dataset performs best with higher physical loss weights. Overall, increasing the $\lambda$ value has a greater impact on datasets with greater variability, while datasets with lower variability (e.g., US-States) maintain good performance even at higher $\lambda$ values. These results suggest that reasonably adjusting the balance between physical and data loss can enhance model performance across different datasets.
IV-G4 Sparse Threshold
To evaluate whether the model is sensitive to the sparsification of the MAG, we test the impact of different sparse thresholds, from 0 to 0.8 (Figures 3 (j), (k), and (l)). A higher threshold corresponds to a sparser MAG. On the Japan-Prefectures dataset, as the sparse threshold increases, RMSE fluctuates slightly, indicating that excessive sparsification may degrade the forecasting performance. On the US-Regions dataset, RMSE shows minor fluctuation, with mild sparsification improving model performance, while PCC remains stable, indicating that the model is not highly sensitive to changes in the sparse threshold. For the US-States dataset, the model achieves the best at a lower sparse threshold, with a slight decrease in PCC as the threshold increases. In summary, a modest increase in the sparse threshold can improve model performance, and selecting an appropriate threshold helps balance data complexity and model effectiveness.
(a) Japan-Prefectures
(b) US-Regions
(c) US-States
(d) Japan-Prefectures
(e) US-Regions
(f) US-States
(g) Japan-Prefectures
(h) US-Regions
(i) US-States
(j) Japan-Prefectures
(k) US-Regions
(l) US-States
@@ IV-G4 Sparse Threshold
Figure 3: Sensitivity analysis on hyperparameter in three real-world datasets. Performance with window size $w$, horizon size $h$, loss weight $\lambda$, and sparse threshold $\delta$. The shaded area represents the 95% confidence interval. Except for the sensitivity analysis experiment of $h$, other sensitivity analysis experiments are conducted under horizon=5.
IV-H Case Study
This section addresses RQ4. To demonstrate the effectiveness of the proposed HeatGNN in capturing spatio-temporal heterogeneity and mechanistic heterogeneity, we conduct a case study to evaluate its short-term and long-term forecasting performance. Specifically, we select two locations from the Japan-Prefectures dataset. We analyze their forecasting performance under the short-term (horizon=2) and long-term (horizon=5) scenarios to explore the forecasting performance of the HeatGNN under different locations. In short-term forecasting (Figures 4(a) and 4(b)), the model accurately predicts the fluctuation trends of epidemic transmission risk. The model’s predictions are close to the actual situation both at the peak and trough of the epidemic, indicating that it is also applicable during extreme events (e.g., 8-10 weeks in Figure 4(a)). This peak sensitivity is critical in the epidemic response strategy, which helps to take interventions in advance to control outbreaks. The model performs stable prediction performance in long-term forecasting, in which the model captures the direction of risk fluctuations (Figures 4(c) and 4(d)). Furthermore, the model can predict a rapid rise in the epidemic 1-2 weeks before an outbreak in long-term forecasting, providing valuable lead time for early prevention efforts. To further validate our findings, we perform a similar analysis using data from the US-States dataset (Figure 5), and the results are consistent with those from the Japan-Prefectures dataset. Those results indicate that HeatGNN effectively captures both short-term and long-term epidemic transmission.
(a) No.0 Location
(b) No.25 Location
(c) No.0 Location
(d) No.25 Location
Figure 4: Case study in Japan-Prefectures. (a) and (b) are conducted under horizon=2. (c) and (d) are conducted under horizon=5.
(a) No.0 Location
(b) No.7 Location
(c) No.0 Location
(d) No.7 Location
Figure 5: Case study in US-Regions. (a) and (b) are conducted under horizon=2. (c) and (d) are conducted under horizon=5.
IV-I Inference Analysis
@@
This section addresses RQ5. To verify the inference efficiency of our framework, we conduct an inference analysis experiment of different location scales based on the US-States dataset to evaluate the inference speed. The results in Figure 6 indicate that the inference time of the HeatGNN remains stable with only minor fluctuations under different location scales. This consistent performance demonstrates the robustness and reliability of HeatGNN, further highlighting its scalability and suitability for applications involving dynamic or large-scale epidemic transmission graph datasets.
@@
V Related Work
V-A Epidemic Forecasting
Epidemic forecasting is a critical task in monitoring the spread of epidemics to help public health interventions quickly and avoid large-scale outbreaks [36]. The time dependence and seasonality of epidemic spread are considered at the beginning of epidemic forecasting. Time series regression models, such as ARIMA[37] and VAR [38], and machine learning models, such as random forests[5], are widely used to solve this task, but these models cannot capture spatial dependencies and ignore the evolution of epidemic transmission. To further improve the ability of the model to capture temporal dependencies, the RNN and RNN-based models were proposed for this task [39]. CNN, Transformer, and their adopted models are used to capture multi-scale features, such as short-term temporal dependencies and long-term temporal dependencies[40]. Meanwhile, the emergence of GNN provides an effective solution to identifying spatio-temporal patterns in epidemic forecasting. In addition to most studies focusing on the impact of geographic topology between locations [4, 14, 12], some studies focus on the distinction between local and global information in epidemic spread [12], as well as multi-scale information [13, 41]. However, these methods mainly focus on the spatio-temporal dependence of epidemic transmission but ignore intrinsic transmission patterns and the evolution of epidemics. They also lack mechanism models to describe the transmission mechanisms behind epidemics. These limitations make it challenging to provide sufficient information to support public health decision-making.
V-B Spatio-temporal Graph Learning
Recently, spatio-temporal graph learning has emerged as a powerful framework for modeling complex data structures that evolve over both spatial and temporal. Graph Convolutional Networks (GCNs) have been developed to achieve end-to-end learning of graph structures[42]. GCNs can efficiently capture local features by sharing parameters in the neighborhood of the graph. Since then, GNNs expanded from static graphs to dynamic graphs to handle the requirement of spatio-temporal data modeling from the needs of dynamic network tasks, like traffic[43], electric load[44], and epidemic[4] forecasting, which involves temporal evolution. The emergence of dynamic graph neural networks, like graph convolutional recurrent networks (GCRN), enables GNNs to update graph structures in time series and capture dynamic changes in node relationships[45]. To further integrate temporal and spatial dependencies, STGNNs combine graph convolution and temporal convolution, enabling the model to capture the dynamic characteristics of temporal and spatial simultaneously, showing significant advantages in various spatio-temporal data tasks [27].
With the development of spatio-temporal graph learning, some studies have introduced recurrent neural networks [46, 47] and Transformer-based attention mechanisms [48] to learn spatio-temporal dependencies more flexibly. In addition, the demand for spatio-temporal dependencies across different scales is driven by the proposed multiscale STGNNs[49]. STGNNs combined with the hypergraph structure are used to process complex high-order interactions in spatio-temporal graph learning[50, 51]. The spatio-temporal graph diffusion network was proposed to describe the information diffusion process in spatio-temporal graph learning [43]. Meanwhile, he heterogeneous STGNNs have been proposed to better adapt to complex heterogeneity problems in spatio-temporal data[52]. Overall, STGNN can capture the spatio-temporal dependencies of epidemics well. Therefore, to enable our proposed framework to capture the spatio-temporal heterogeneity of various epidemic locations, we are inspired by [12] and use it as the spatio-temporal graph module.
V-C Physics-guided Machine Learning
The approximations learned by machine learning methods often ignore or violate existing physical knowledge, leading to misleading or unexplainable conclusions [53]. Therefore, the combination of physics and machine learning has begun to be proposed to solve the problem of lack of interpretability and black-box structures in machine learning. A commonly used method is physics-informed neural networks (PINN) [28], which is based on combining physical residual loss with data loss to transfer or distill physical epidemic knowledge into a machine learning framework for constrained learning processes while being able to effectively capture heterogeneous data in different forecasting tasks, such as urban flow[54], blood pressure[55], rumor detection[56], and epidemic transmission[57, 19]. Another way to combine physical models and machine learning is Neural differential equations (Neural ODEs), which use the neural network as a continuous dynamic process. For example, the time-dependent states in the epidemic model are replaced by neural networks[58]. Neural ODEs are also used in fields such as climate, weather[59], polar motion[60], and battery health[61] forecasting to solve real-world complex continuous-time systems.
Some scenarios in epidemic transmission require integrating physical models into machine learning models to improve performance and interpretability. Epidemic models are based on overly simple assumptions and cannot fully understand the complex epidemic spread[9], while machine learning models lack effective mechanism explanations. Extended epidemic models with networks have been proposed to describe the transmission patterns between locations[62, 11]. However, these network-based epidemic models have problems with large-scale parameter optimization[11] or lack of data[63, 21]. To address those problems, some models proposed to use simulated data generated by physical models to solve the problem of insufficient data and provide high-resolution forecasting[17, 63]. Meanwhile, machine learning has also begun to estimate epidemic models to solve the parameter explosion [21, 20]. Neural ODEs are incorporated into epidemic models to learn the continuous evolution of epidemic spread in different regions [18]. The epidemic models were used as causal mechanism models to guide the graph embedding learning process to handle noisy data and missing data better[64]. Although physics-guided machine learning has made efforts in epidemic forecasting, previous studies have ignored the mutual influence and dependency of the intrinsic epidemic transmission between locations, which makes it challenging to forecast epidemics breaking out in multiple locations.

### v2

@@ V-D Ablation Study (RQ2)
We conduct the ablation study by testing the following variants of HeatGNN: (1) “w/o PL”, remove the physics loss from HeatGNN, i.g. remove the constraints on the EIEL module; (2) “w/o TG”, remove the TG module and use only the STGL module for forecasting while retaining the EIEL module to guide the output of the framework; (3) “w/o TG+EIEL”, remove the EIEL and TG modules and only keep the STGL module for forecasting due to the TG module being based on the EIEL module.
@@
The ablation study results of RMSE and PCC are shown in Table IV. The results show that our proposed HeatGNN performs robustly across most cases. The physical loss incorporates domain knowledge of epidemic transmission, particularly benefiting high-variability regions (e.g., Japan-Prefectures and US-Regions) and long-term forecasting. Removing the physics loss (“w/o PL”) results in the model lacking intrinsic epidemic transmission mechanism constraints, which increases the forecasting error and reduces its ability to capture trends. However, its impact is limited in the US-States dataset, which has low volatility and relatively uniform data. In fact, the “w/o PL” variant performs comparably to HeatGNN on the US-State dataset at some forecast horizons, suggesting that the benefits of physical loss’s regularization are less significant in more stable, homogeneous regions.
The TG module is crucial for modeling cross-regional epidemic transmission. After removing the TG module (“w/o TG”), the model is significantly weaker at capturing the intrinsic epidemic transmission relationship between locations. While it has a similar role to the physics loss across datasets, its impact on performance is more pronounced. This is particularly evident in the US-Region and US-State datasets, where the “w/o TG” leads to a significant drop in PCC and an increase in RMSE. Moreover, on the Australia-COVID dataset, “w/o TG” slightly improves RMSE at longer horizons, suggesting that TG may introduce redundancy when inter-regional transmission is weak, while PCC still drops, indicating that TG remains useful for capturing trend consistency.
Furthermore, the EIEL module enhances the effectiveness of the TG module, further improving the ability to capture epidemic transmission. Relying solely on the STGL module is insufficient for addressing complex epidemic forecasting tasks. When both TG and EIEL modules are removed (“w/o TG+EIEL”), the model struggles to accurately capture epidemic transmission over extended time horizons, resulting in a significant decline in forecasting performance. These findings highlight the complementary and synergistic roles of each module. The complete HeatGNN consistently achieves the best result in most cases and forecast horizons, demonstrating the importance of each component and its integration.
TABLE IV: Ablation test results on four datasets: Japan-Prefectures, US-Region, US-State, and Australia-COVID.
Method Japan-Prefectures US-Region
RMSE ($\downarrow$) PCC ($\uparrow$) RMSE ($\downarrow$) PCC ($\uparrow$)
2 5 7 12 2 5 7 12 2 5 7 12 2 5 7 12
w/o PL 1.230 1.453 1.804 1.759 0.900 0.863 0.746 0.757 0.631 0.943 1.055 1.141 0.924 0.864 0.840 0.823
w/o TG 1.229 1.439 1.762 1.657 0.891 0.871 0.783 0.754 0.740 1.039 1.142 1.182 0.923 0.838 0.797 0.780
w/o TG+EIEL 1.279 1.452 1.765 1.675 0.890 0.853 0.774 0.760 0.727 0.977 1.119 1.188 0.927 0.854 0.818 0.790
HeatGNN 1.149 1.378 1.735 1.685 0.917 0.884 0.780 0.773 0.541 0.852 0.922 1.024 0.941 0.866 0.856 0.824
Method US-State Australia-COVID
RMSE ($\downarrow$) PCC ($\uparrow$) RMSE ($\downarrow$) PCC ($\uparrow$)
2 5 7 12 2 5 7 12 2 5 7 12 2 5 7 12
w/o PL 0.146 0.188 0.192 0.223 0.951 0.917 0.912 0.879 0.487 0.460 0.451 0.455 0.986 0.985 0.985 0.987
w/o TG 0.149 0.202 0.215 0.238 0.951 0.907 0.891 0.858 0.350 0.366 0.362 0.429 0.991 0.991 0.991 0.986
w/o TG+EIEL 0.153 0.206 0.217 0.227 0.951 0.902 0.892 0.880 0.461 0.481 0.449 0.435 0.984 0.986 0.988 0.987
HeatGNN 0.142 0.186 0.200 0.240 0.953 0.921 0.911 0.870 0.315 0.334 0.411 0.466 0.993 0.993 0.991 0.986
@@
Figure 5: Sensitivity analysis on hyperparameter. Performance with window size $w$, horizon size $h$, loss weight $\lambda$, and sparsity threshold $\delta$. Except for the sensitivity analysis of $h$, other sensitivity analyses are conducted under horizon=5.
V-E Hyperparameter Sensitivity Analysis (RQ3)
We test how performance varies with hyperparameters.
Window size $w$. To evaluate whether the model is sensitive to the length of historical data, we evaluate different window sizes ranging from 20 to 60. Figures 5 (a), (b), (c), and (d) show the RMSE and PCC results in different window sizes. For the Japan-Prefectures and Australia-COVID datasets, forecasting performance improves significantly with larger window sizes, as evidenced by a decrease in RMSE and an increase in PCC. In contrast, model performance remains stable across all window sizes for the US-Regions dataset, which indicates that the model is not sensitive to the window size change in this dataset. For the US-States dataset, the optimal performance is observed at a window size of 60. In summary, while increasing the window size enhances model performance, the improvement becomes limited after a threshold.
Horizon size $h$. To evaluate the sensitivity of the model to the size of the forecasting horizon, we analyze its performance across horizons ranging from 1 to 6. Figures 5 (e), (f), (g), and (h) show the RMSE and PCC results for different horizons. In the Japan-Prefectures dataset, as the horizon increases, RMSE gradually rises while PCC decreases. Similar trends are observed in the US-Regions, US-States, and Australia-COVID datasets. Comparing the results across the four datasets, we find that datasets with more significant fluctuations (Japan-Prefectures), shorter horizons, and more accurate predictions, while datasets with lower variability (US-Regions, US-States, and Australia-COVID) maintain robust forecasting performance even with longer horizons.
Loss weight $\lambda$. To evaluate the sensitivity of the model to the balance weight between physical and data loss, we examine the effect of varying $\lambda$ values from 0 to 1 on forecasting performance. Figures 5 (i), (j), (k), and (l) show the RMSE and PCC results under different $\lambda$. Overall, increasing the $\lambda$ value has a greater impact on datasets with greater variability, while datasets with lower variability (e.g., US-States) maintain good performance even at higher $\lambda$ values. The results suggest that reasonably adjusting the balance between physical and data loss can enhance model performance across different datasets.
Sparsity Threshold $\delta$. To evaluate whether the model is sensitive to the sparsification of the MAG, we test the impact of different sparsity thresholds, from 0 to 0.8 (Figures 5 (m), (n), (o), and (p)). A higher threshold corresponds to a sparser MAG. On the Japan-Prefectures dataset, as the sparsity threshold increases, RMSE fluctuates, indicating that excessive sparsification may degrade the forecasting performance. On the US-Regions dataset, RMSE shows minor fluctuation, while PCC remains stable. For the US-States dataset, the model achieves the best performance at a lower sparsity threshold, with a slight decrease in PCC as the threshold increases. For the Australia-COVID dataset, the performance of different sparsity thresholds basically remains unchanged. Overall, a modest increase in the sparsity threshold improves model performance, and selecting an appropriate threshold helps balance data complexity and model effectiveness.
V-F Interpretability Analysis (RQ4)
We conduct a case study in the Japan-Prefectures dataset with horizon=3 and 47 nodes. Figures 6 and 7 illustrate the mechanistic similarity $\boldsymbol{M}_{t}$ between different locations learned from the TG module in the Japan-Prefectures dataset at single and continuous timestep. In this analysis, we focus on week 54 and utilize historical data from week 54 and the previous 19 weeks (a total of 20 weeks) to predict the infection number three weeks later (week 57). By observing location 19 (prediction target), location 36, and location 37, the infection number trends of these three locations show similar patterns up to week 54. However, after week 54, the trend at location 37 diverges significantly, rising sharply in week 57 and reaching a higher peak in subsequent weeks. In contrast, locations 19 and 36 exhibit more similar trends, with a gradual increase in infection number that does not reach the same peak as location 37. The mechanism similarity, calculated using HeatGNN, highlights differences in future mechanism propagation across locations as of week 54. Specifically, the mechanism similarity reveals a lower similarity between location 19 and location 37 (indicated in blue) and a higher similarity between location 19 and location 36 (indicated in red). This finding demonstrates that the mechanism similarity effectively captures hidden dynamics between locations under the mechanism of spread.
We conduct a MAG comparison experiment to prove that the MAG used by the TG module better captures mechanistic heterogeneity: based on real data, we fit the SIR model for each location and construct a real MAG that reflects the heterogeneity of the real epidemic transmission mechanistic as a baseline (real MAG). Comparing this real MAG with the MAG generated by the EIEL module embedding ($MAG_{E}$) and the MAG generated by the ST module embedding ($MAG_{S}$) based on the Manhattan distance and the Euclidean distance, the results show that the MAG generated by the EIEL module is closer to the real MAG (Table V). This finding indicates that more correlation between EIEL module embedding and established epidemic factors. Furthermore, we calculate the average degree and degree distribution at different thresholds in MAG (Table VI and Figure 8). The motivation behind MAG is to emphasize information propagation between mechanistically similar locations, which is enabled primarily by the similarity-weighted nature of edges. Meanwhile, the role of the threshold is mainly for efficient graph convolution by allowing MAG to be operated as a sparse matrix, thus having a lower impact on the performance. This is also verified by aligning Figure 5 and Table IV’s results - as long as MAG is in use (despite the value of $\delta$), the performance outperforms the “w/o TG” variant (MAG is completely disabled) in most cases.
Figure 6: Mechanistic similarity $\boldsymbol{M}_{t}$. On the left is the actual infection number change of three locations. On the right is the mechanism similarity matrix learned from historical data (up to week 54) for prediction in week 57.
Figure 7: Mechanistic similarity $\boldsymbol{M}_{t}$ in continuous timesteps.
TABLE V: Distances between the generated and real MAGs.
$MAG_{E}$ $MAG_{S}$
Manhattan distance $(\downarrow)$ 9.94 7.29
Euclidean distance $(\downarrow)$ 0.30 0.22
TABLE VI: Average degrees at different thresholds in MAG.
Threshold $\delta$ 0 0.2 0.4 0.6 0.8
Average Degree 47.0 45.7 45.1 43.0 36.2
Figure 8: The degree distribution at different thresholds in MAG.
V-G Time Complexity and Scalability Analysis (RQ5)
This section aims to test whether HeatGNN has robust scalability. We analyze the time complexity of the MAG in HeatGNN. Let $N$ be the number of positions and $N_{f}$ be the feature dimension of each position. The time complexity of MAG includes three parts: the complexity of feature normalization is $O(N\cdot N_{f})$, which is used to standardize the features of each position; the calculation of the cosine similarity matrix in MAG requires similarity evaluation of all position pairs, with a complexity of $O(N^{2}\cdot N_{f})$; and the final MAG normalization operation is $O(N^{2})$. In summary, the overall time complexity of MAG is $O(N^{2}\cdot N_{f})$, among which the cosine similarity calculation is dominant. We compare the inference time of HeatGNN with two baselines (Table VII). The result shows that HeatGNN performs well in inference efficiency, which is lower than other baselines in most cases, indicating that it has good practical application insight while maintaining low computational overhead.
@@ V-G Time Complexity and Scalability Analysis (RQ5)
To verify the inference efficiency of HeatGNN, we conduct an inference analysis of different location scales based on the US-States dataset to evaluate the inference time. The results in Figure 9 indicate that the inference time of the HeatGNN remains stable with only minor fluctuations under different location scales. This consistent performance demonstrates the robustness and reliability of HeatGNN, further highlighting its scalability and suitability for applications involving dynamic or large-scale epidemic transmission graph datasets.
@@
TABLE VII: Comparison of inference time (in second) of different models on four datasets with horizon=5.
Japan-Prefectures US-Region US-State Australia-COVID
Cola-GNN 0.650 0.360 0.720 0.735
Epi-Cola-GNN 1.029 1.108 1.060 0.236
HeatGNN 0.370 0.231 0.241 0.265
V-H Robustness Analysis (RQ6)
Unlike purely data-driven methods, HeatGNN integrates an epidemic model and uses PINN to embed epidemic dynamics into learning. Ensuring consistency with known transmission mechanisms reduces reliance on noisy or incomplete data, improving robustness. We further test with different noisy node proportions by adding Gaussian noise and different missing node proportions by removing nodes, where the results are shown in Table VIII. The results indicate that HeatGNN outperforms Cola-GNN and EpiGNN in most cases under different noise and missing node ratios. The results indicate HeatGNN’s robustness and prediction stability under incomplete graph structures, and it can more effectively support epidemic transmission prediction tasks in complex public health scenarios.
TABLE VIII: Performance under different noisy and missing node proportions in the US-Region with horizon=2.
RMSE $\times 10^{3}$ ($\downarrow$) PCC ($\uparrow$)
Perturbation Proportion 0% 10% 20% 30% 0% 10% 20% 30%
Noisy Cola-GNN 0.552 0.572 0.601 0.615 0.939 0.933 0.925 0.920
EpiGNN 0.563 0.584 0.619 0.634 0.937 0.930 0.920 0.915
HeatGNN 0.541 0.561 0.597 0.615 0.941 0.935 0.926 0.920
Missing Cola-GNN 0.552 0.906 1.155 1.326 0.939 0.827 0.726 0.648
EpiGNN 0.563 0.947 1.167 1.327 0.937 0.808 0.714 0.636
HeatGNN 0.541 0.906 1.094 1.247 0.941 0.822 0.742 0.669